### v5

# A Codimension Two Approach to The $\mathbb{S}^{1}$-Stability Conjecture

Abstract.
J. Rosenberg’s ${\mathbb{S}}^{1}$-stability conjecture states that a closed oriented manifold $X$ admits a positive scalar curvature metric iff $X\times{\mathbb{S}}^{1}$ admits a positive scalar curvature metric $h$. As pointed out by J. Rosenberg and others, there are known counterexamples in dimension four. Using conformal geometry techniques, we prove this conjecture whenever $h$ satisfies a geometric bound which measures the discrepancy between $\partial_{\theta}\in T{\mathbb{S}}^{1}$ and the normal vector field to $X\times\{P\}$, for a fixed $P\in\mathbb{S}^{1}.$
1. Introduction
In [8, Conj. 1.24], J. Rosenberg proposed the $\mathbb{S}^{1}$-stability conjecture for metrics of positive scalar curvature (PSC).
${\mathbb{S}}^{1}$-Stability Conjecture. Let $X$ be a closed oriented manifold. Then $X$ admits a PSC metric if and only if $X\times\mathbb{S}^{1}$ admits a PSC metric.
In [8, Rmk. 1.25], Rosenberg gave a counterexample in dimension four of an odd degree hypersurface in $\mathbb{C}\mathbb{P}^{3}$, based on Seiberg-Witten theory. Other counterexamples in dimension four are in [6, Rmk. 5]. The conjecture has been verified in dimensions $2,3,5,6$ and for a large class of spin manifolds [10, Cor. 1.5], using a combination of minimal surface methods and band-width estimates. (See e.g., [3, §1.1] for the ingredients for dimension 2, [5, Cor. 7.34] for dimension 3, and [7, Rmk. 2.26] for dimensions 5, 6.)
The nontrivial direction of the conjecture states that if $(X\times{\mathbb{S}}^{1},h)$ is a PSC metric, then $X$ admits a PSC metric. We introduce a PDE approach to solve the $\mathbb{S}^{1}$-stability conjecture in all dimensions under a restriction on $h$. To state the main theorem, let $\partial_{\theta}$ be the usual tangent vector field to ${\mathbb{S}}^{1}$, and let $\mu$ be the unit normal vector field to the slice $X\times\{P\}$ for a fixed $P\in{\mathbb{S}}^{1}$, where $\mu$ is chosen so that $h(\mu,\partial_{\theta})>0.$ Define the $h$-angle $\angle_{h}(\mu,\partial_{\theta})$ by $h(\mu,\partial_{\theta})=\|\mu\|_{h}\|\partial_{\theta}\|_{h}\cos(\angle_{h}(\mu,\partial_{\theta}))=\|\partial_{\theta}\|_{h}\cos(\angle_{h}(\mu,\partial_{\theta}))$.
Theorem 3.1.
Let $X$ be an oriented closed manifold with $\dim X\geqslant 2$. If $X\times\mathbb{S}^{1}$ admits a PSC metric $h$ such that for some $P\in{\mathbb{S}}^{1}$, we have
(1) $\angle_{h}(\mu,\partial_{\theta})<\frac{\pi}{4}$
on $X\times\{P\},$ then $X\approx X\times\{P\}$ admits a PSC metric.
To be precise, the ${\mathbb{S}}^{1}$-conjecture holds on $X\times\{P\}$ if $\partial_{\mu}$ lies inside the cone with axis $\partial_{\theta}$ and $h$-angle $\pi/4$ on each $T_{(x,P)}(X\times{\mathbb{S}}^{1})$, for all $x\in X.$
The main techniques used are a mix of the Gauss-Codazzi equation and conformal geometry, along with the introduction of an extra dimension. To outline the proof of the main theorem, we refer to the diagram (2). We take the PSC metric $h$ on $X\times{\mathbb{S}}^{1}$ and extend it to a product metric $g$ on $M:=W\times{\mathbb{S}}^{1}:=X\times[-1,1]\times{\mathbb{S}}^{1}$. Note that $X$ has codimension two in $M$. This induces a metric $\sigma^{*}g$ on $W$, which we use to construct a PDE (7), which is elliptic if (1) holds. This PDE is related to a conformal transformation of the Gauss-Codazzi equation, and the choice of the inhomogeneous term $F$ in (7) is crucial. Using the ellipticity, we find a smooth solution $u$ to (7) on $W$, such that $u_{W}:=u+1$ is positive. We pull back $u_{W}$ to $u_{M}$ on $M$, and use $u_{M}$ to define a conformal transformation of $g$ to $\tilde{g}=u_{M}^{\frac{4}{n-2}}g$. Restricting $\tilde{g}$ to $X\times S^{1}$ and then to $X$ gives a PSC metric $\tau^{*}\imath^{*}\tilde{g}$ on $X$.
It is natural to ask why the extra dimension $[-1,1]$ is introduced. If we take a conformal transformation $\tilde{h}=u^{\frac{4}{n-2}}h$ on $X\times S^{1}$ and use the Gauss-Codazzi equation to try to prove that the scalar curvature of $X\times\{P\}$ is positive, we find that $u$ must satisfy a non-elliptic PDE. In particular, we cannot guarantee the existence of a smooth solution with the necessary small $\mathcal{C}^{1,\alpha}$ estimate (see Prop. 2.1). By adding in the extra dimension, we produce and solve (7), provided (1) holds. This is discussed before Lem. 2.1.
As an outline of the paper, in §2 we prove a series of technical results. The elliptic PDE (7) on $W$ related to the conformal transformation $g\mapsto\tilde{g}$ is given in Prop. 2.1. We prove that (7) has a smooth solution $u$ with $\|u\|_{\mathcal{C}^{1,\alpha}}$ arbitrarily small, by choosing the inhomogeneous term $F$ to be concentrated near $X=X\times\{0\}.$ In particular, Lem. 2.3 gives a partial $\mathcal{C}^{2}$-estimate for $u_{W}$, where the hypothesis that $h$ is PSC is essentially used. We also prove several technical lemmas which allow us to compare the Laplacians and gradients on $M,W,$ and $X\times{\mathbb{S}}^{1}$.
In §3 we prove the main theorem Thm.  3.1, essentially by using Gauss-Codazzi on $\tilde{g}$ twice to move from $M$ back to $X\times{\mathbb{S}}^{1}$ and then back to $X$. The technical results from §2 appear in the Gauss-Codazzi formulas and various conformal transformation rules in the proof of Thm.  3.1.
We thank B. Hanke, D. Ruberman, T. Schick and B. Sen for corrections to an earlier version of this paper.
2. Technical Preliminaries
In this section, we prove preliminary results that are used in the proof of Thm.  3.1 in §3.
We introduce the setup and notation. $X$ is an oriented closed manifold such that $X\times\mathbb{S}^{1}$ has a PSC metric $h$. We assume dim$(M)=n-1\geq 2.$ Set
$Y=X\times{\mathbb{S}}^{1},\ W=X\times[-1,1],\ M=W\times\mathbb{S}^{1}=Y\times[-1,1].$
$X\times\mathbb{S}^{1}=X\times\{0\}\times{\mathbb{S}}^{1}$ is an embedded hypersurface of $M$.
Put the product metric $g=h\oplus dt^{2}$ on $M$, for $t\in[-1,1].$ For a fixed $P\in\mathbb{S}^{1}$, define $\sigma:W\rightarrow W\times\mathbb{S}^{1}$ by $\sigma(w)=(w,P),$ and $\tau:X\rightarrow X\times\mathbb{S}^{1}$, $\tau(x)=(x,P).$ For the inclusion maps $\imath=\imath_{0}:X\rightarrow W$ and $\imath=\imath_{0}:X\times\mathbb{S}^{1}\rightarrow M$, where $\imath_{0}(x)=(x,0)$ and $\imath_{0}(x,P)=(x,0,P)$, we have induced metrics $\sigma^{*}g$ on $W$ and $\tau^{*}\imath^{*}g=\tau^{*}h$ on $X$. The following diagram summarizes the setup.
(2) ${(Y=X\times{\mathbb{S}}^{1},h=\imath^{*}g)}$${(M=W\times{\mathbb{S}}^{1}=Y\times[-1,1],g)}$${(X,\tau^{*}\imath^{*}g=\imath^{*}\sigma^{*}g)}$${(W=X\times[-1,1],\sigma^{*}g)}$$\scriptstyle{\imath}$$\scriptstyle{\imath}$$\scriptstyle{\tau}$$\scriptstyle{\sigma}$
We will frequently identify $x\in X$ with $(x,0)\in W$, $y:=(x,P)\in Y$, and $(x,0,P)\in M.$ Note that dim $M\geq 4.$
Let $\theta$ be the usual coordinate on ${\mathbb{S}}^{1}$ with corresponding global vector field $\partial_{\theta}$ on $W$ and $M$. Let $\mu$ be the outward normal vector field to $X\times\{Q\}$, for all $Q\in{\mathbb{S}}^{1},$ chosen so that $h(\mu,\partial_{\theta})>0.$ $\mu$ is a global vector field on $W$. Locally on $W$, we can consider $\mu$ to be the tangent vector $\partial_{x^{n}}$, where $(x^{1},\ldots,x^{n-1})$ are local coordinates on $U\subset X$. In these coordinates, $h(\mu,\mu)=h_{nn}=g_{nn}:=g_{\mu\mu}$, where the expressions in $g$ are valid on $U\times[-1,1].$ Below, we will drop some inclusion maps and write $g_{nn}$ on both $M$ and $X\times\{0\}\times\mathbb{S}^{1}.$
As outlined in the Introduction, in this section we solve a PDE (7) involving $\sigma^{*}g$ on $W$ (Prop. 2.1). This PDE is elliptic under the condition (1) (Lem.  2.1). We modify a solution $u$ to a positive function $u_{W}$ on $W$, pull back $u_{W}$ to $u_{M}$ on $M$, and use $u_{M}$ to conformally change $g$ to $\tilde{g}=u_{M}^{\frac{4}{n-2}}g$. The main theorem in §3 proves that for a careful choice of the inhomogeneous term $F$ in (7), $\tau^{*}\imath^{*}\tilde{g}$ is PSC, provided (1) holds. To this end, the diagram indicates that we have to compare Laplacians $\Delta_{g}$, $\Delta_{\imath^{*}g}$ and $\Delta_{\sigma^{*}g}$ (Lem. 2.4), the scalar curvatures of $\tilde{g}$ and $\imath^{*}\tilde{g}$, and the scalar curvatures of $\imath^{*}\tilde{g}$ and $\tau^{*}\imath^{*}\tilde{g}$ (see the proof of Thm.  3.1) via the Gauss-Codazzi equation.
The function $u_{M}$ defined above is independent of $\theta.$ However, comparing the scalar curvatures for induced metrics via Gauss-Codazzi requires the use of the normal vector $\mu.$ For the product metric on $X\times S^{1}$, $\mu=\partial_{\theta}$, so in the non-product case we have to compare these vector fields. On $X\times S^{1}$, set
(3) $\mu=a\partial_{\theta}+\sum_{i=1}^{n-1}b^{i}\partial_{i}:=a\partial_{\theta}+V,$
for a local coordinate frame $\{\partial_{i}\}$ on $X$. Note that
$1=h(\mu,\mu)=h(a\partial_{\theta},\mu)+b^{i}h(\partial_{i},\mu)=ah(\partial_{\theta},\mu)\Rightarrow a=h(\partial_{\theta},\mu)^{-1},$
where clearly $h(\partial_{\theta},\mu)\neq 0.$ Since $\mu,a\partial_{\theta}$ are globally defined vector fields, so is the vector field $V$ on $X\times\{P\}$ for each $P\in{\mathbb{S}}^{1}.$
Finally, $\mu,\partial_{\theta}$ and hence $V$ pull back to $t$-independent vector fields on $M$. Since $\sigma=\sigma_{P}$ identifies $X\times[-1,1]\times\{P\}$ with $W$, it is easy to check that $|V_{(x,P)}|^{2}_{h}=|V_{(x,t)}|^{2}_{\sigma^{*}g}$ for all $t$.
The operator $L^{\prime}$ in the next lemma appears in the proof of the main theorem; see (3). This lemma motivates the introduction of the extra dimension $[0,1]$ in $W$; if we work on $X\times{\mathbb{S}}^{1}$, the operator analogous to $L^{\prime}$ associated to a conformal change of $h$ would be $\nabla_{\mu}\nabla_{\mu}-\Delta_{h}$, which is never elliptic.
Note that (4) is equivalent to the angle hypothesis (1) in Thm.  3.1.
Lemma 2.1.
Fix $P\in{\mathbb{S}}^{1},$ and identify $X\times[-1,1]\times\{P\}$ with $W$. If
(4) $\frac{h(\partial_{\theta},\partial_{\theta})}{h(\mu,\partial_{\theta})^{2}}<2,$
then the operator
$L^{\prime}:=\nabla_{V}\nabla_{V}-\Delta_{\sigma^{*}g}$
is elliptic on $W$.
We use the convention that $-\Delta_{\ell}$ is positive definite for a metric $\ell.$
Proof.
In local coordinates,
$\nabla_{V}\nabla_{V}=\sum_{i,j=1}^{n-1}b^{i}(x)b^{j}(x)\frac{\partial^{2}}{\partial x^{i}\partial x^{j}}+G_{1}(x_{0})$
where $G_{1}(x)$ is a linear first order operator. In Riemannian normal coordinates $(x^{1},\ldots,x^{n-1})$ centered at a fixed $x_{0}\in X$, we have
$\Delta_{\sigma^{*}g}|_{x_{0}}=\sum_{i=1}^{n-1}\frac{\partial^{2}}{\partial\left(x^{i}\right)^{2}}+\frac{\partial^{2}}{\partial t^{2}}.$
The principal symbol of $L^{\prime}$ at $x_{0}$ is
$\sigma_{2}(L^{\prime})(x_{0},\xi)=-\sum_{i,j=1}^{n-1}b^{i}(x_{0})b^{j}(x_{0})\xi^{i}\xi^{j}+\sum_{i=1}^{n-1}\left(\xi^{i}\right)^{2}+(\xi^{n})^{2}.$
Proving $\sigma_{2}(L^{\prime})(x_{0},\xi)>0$ for $\xi\neq 0$ is equivalent to proving that
$B:=-\begin{pmatrix}(b^{1})^{2}&b^{1}b^{2}&\dotso&b^{1}b^{n-1}\\ b^{2}b^{1}&(b^{2})^{2}&\dotso&b^{2}b^{n-1}\\ \vdots&\vdots&\ddots&\vdots\\ b^{n-1}b^{1}&b^{n-1}b^{2}&\dotso&(b^{n-1})^{2}\end{pmatrix}+I_{n-1}:=B^{\prime}+I_{n-1}$
is positive definite. Sylvester’s criterion states the symmetric matrix $B$ is positive definite iff all its $k\times k$ principal minors $B_{k}$ have positive determinants.
We claim that
(5) $\det(B_{k})=1+{\rm Tr}(B_{k}^{\prime})=1-\sum_{i=1}^{k}(b^{i})^{2},k=1,\dotso,n-1.$
We show this for $B_{n-1}=B$, as the argument for the other $B_{k}$ is identical. We first observe that if $b^{i}\neq 0,\forall i,$ then all rows of $B^{\prime}$ are proportional to each other. It follows that the kernel of $B^{\prime}$ has dimension $n-2$, so 00 is an eigenvalue of $B^{\prime}$ of multiplicity $n-2$. Since the sum of the eigenvalues of $B^{\prime}$ is the trace $-\sum_{i=1}^{n-1}(b^{i})^{2}$, the nontrivial eigenvalue must be $-\sum_{i=1}^{n-1}(b^{i})^{2}$. Thus the eigenvalues of $B$ are
$\lambda_{1}=-\sum_{i=1}^{n-1}(b^{i})^{2}+1,\lambda_{2}=\ldots=\lambda_{n-1}=1,$
which proves (5). If some $b^{i}$ vanish, then the claim reduces to a lower dimensional case by removing the corresponding rows and columns of all zeros.
Therefore, $L^{\prime}$ is elliptic if
(6) $\sum_{i=1}^{n-1}(b^{i})^{2}<1.$
In normal coordinates,
$\sum_{i=1}^{n-1}(b^{i})^{2}=|V|^{2}_{\sigma^{*}g}=|V|^{2}_{h}=h(\mu-h\left(\mu,\partial_{\theta})^{-1}\partial_{\theta},\mu-h(\mu,\partial_{\theta})^{-1}\partial_{\theta}\right)=1-2+\frac{h(\partial_{\theta},\partial_{\theta})}{h(\mu,\partial_{\theta})^{2}}.$
Thus (6) holds if
$\frac{h(\partial_{\theta},\partial_{\theta})}{h(\mu,\partial_{\theta})^{2}}<2.$
Using this lemma, we will construct an elliptic PDE with an appropriate inhomogeneous term $F$ whose solution has small $\mathcal{C}^{1,\alpha}$-norm. A solution of this PDE plays a crucial role in the proof of Thm.  3.1. We begin by finding a function $F$ that is sufficiently large on $X\times\{0\}$, but with small $\mathcal{L}^{p}$-norm on $W$.
Lemma 2.2.
Fix $p\in\mathbb{N}$, $C\gg 1$, and a positive $\delta\ll 1$. For any positive $\epsilon<1,$ there exists a positive smooth function $F=F_{\epsilon}:W\to\mathbb{R}$ such that $F|_{X\times[-\epsilon/2,\epsilon/2]}=C+1$ and $\lVert F\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}<\delta.$
Proof.
Given $p,C,\delta$, set $f=C+1$ on $X$. For a positive constant $\epsilon\ll 1$, construct a positive smooth function $\phi:[-1,1]\rightarrow[0,1]$, with $\phi(t)=1$ for $|t|\in\left[0,\frac{\epsilon}{2}\right],$ and $\phi(t)=0$ for $\lvert t\rvert\in[\epsilon,1]$. Set $F=f\cdot\phi:W\rightarrow\mathbb{R}.$ Then $F=C+1$ on $X{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times[-\epsilon/2,\epsilon/2]}$ and $F>0$ on $W$. Finally,
$\|F\|^{p}_{\mathcal{L}^{p}(W,\sigma^{*}g)}=\int_{X\times[-\epsilon,\epsilon]}|F|^{p}\text{dvol}_{\sigma^{*}g}<\delta,$
for $\epsilon$ sufficiently small. ∎
Let $\nu=\pm\partial_{t}$ be the unit outward normal vector field along $\partial W$, and let $R_{\ell}$ be the scalar curvature of a metric $\ell$. Define the $\mathcal{C}^{1,\alpha}$ norm of $u:W\to\mathbb{R}$ by choosing a cover $\{(U_{i},(x^{k}_{i}))\}$ of $W$ and setting
$\|u\|_{\mathcal{C}^{1,\alpha}}=\|u\|_{\mathcal{C}^{0,\alpha}}+\sup_{i,k}\|\partial_{x_{i}^{k}}u\|_{\mathcal{C}^{0,\alpha}}.$
Proposition 2.1.
Let $(W,\sigma^{*}g)$ be as above. Assume that (1) holds. For any positive constant $\eta\ll 1$, there exist constants $p,\delta,C,\epsilon$ and an associated $F$, in the sense of Lem. 2.2, such that for fixed $P\in\mathbb{S}^{1}$,
(7) $4\nabla_{V}\nabla_{V}u-4\Delta_{\sigma^{*}g}u+R_{g}|_{\sigma(W)}u=F\;{\rm in}\;W,\ u=0\;{\rm on}\;\partial W,$
admits a smooth solution $u:W\to\mathbb{R}$ with
(8) $\lVert u\rVert_{\mathcal{C}^{1,\alpha}(W)}\leqslant\eta.$
In the proof and the rest of the paper, norms of gradients like $|\nabla_{\sigma^{*}g}u|_{\sigma^{*}g}^{2}$ are usually just denoted by $|\nabla_{\sigma^{*}g}u|^{2}.$
Proof.
Denote the $\mathcal{L}^{p}$ Sobolev spaces by $\mathcal{W}^{k,p}.$ Fix $p\gg 0$ so that the Sobolev embedding
$\mathcal{W}^{2,p}(W,\sigma^{*}g)\hookrightarrow\mathcal{C}^{1,\alpha}(W)$ is a compact inclusion. Define the elliptic operator
$L:=4\nabla_{V}\nabla_{V}u-4\Delta_{\sigma^{*}g}+R_{g}|_{\sigma(W)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=4L^{\prime}+R_{g}|_{\sigma(W)}}$
on $W$. The Dirichlet boundary condition allows us to rewrite (7) on $W$ in the weak form
${-\int_{W}4\nabla_{V}u\cdot\nabla_{V}\phi\ \text{dvol}_{\sigma^{*}g}+\int_{W}4\nabla_{\sigma^{*}g}u\cdot\nabla_{\sigma^{*}g}\phi\ \text{dvol}_{\sigma^{*}g}+\int_{W}uL_{0}(\phi)\ \text{dvol}_{\sigma^{*}g}}$
(9) $\qquad+\int_{W}R_{g}|_{\sigma(W)}u\phi\ \text{dvol}_{\sigma^{*}g}$
$=\int_{W}F\phi\ \text{dvol}_{\sigma^{*}g},\forall\phi\in\mathcal{C}_{c}^{\infty}(W),$
where $L_{0}$ is a first order operator determined by $V$. Note that
(10) $R_{h}>0\ \text{on}\ X\times{\mathbb{S}}^{1}\Rightarrow R_{g}>0\ \text{on}\ M.$
The operator $L$ is injective: if $Lu=0$, then combining Lem. 2.1, (10) and the maximum principle gives $u=0.$ It follows from Fredholm alternative that (2) has a weak solution in $\mathcal{W}^{1,2}(W)$. A standard bootstrapping argument [4, Thm.  1] then implies $u$ is smooth. By classical elliptic regularity theory on $(W,\partial W,\sigma^{*}g)$ with Dirichlet boundary condition [1], there exists a constant $C_{1}=C_{1}(W,\sigma^{*}g,p)$ such that
(11) $\lVert u\rVert_{\mathcal{W}^{2,p}(W,\sigma^{*}g)}\leqslant C_{1}\left(\lVert Lu\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}+\lVert u\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}\right),$
for $u\in\mathcal{L}^{p}(W,\sigma^{*}g)\cap\mathcal{W}^{1,2}(W,\sigma^{*}g)$.
We claim that for all $u\in\mathcal{W}^{1,2}(W)\cap\mathcal{L}^{p}(W)$ solving (2) weakly, there exists $C_{1}^{\prime}>0$ such that
(12) $\lVert u\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}\leqslant C_{1}^{\prime}\lVert Lu\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}.$
If not, there exists a sequence $\{u_{n}\}$ solving (2) weakly such that
$\lVert u_{n}\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}=1,u_{n}|_{\partial W}=0\;{\rm weakly},\lVert Lu_{n}\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}\leqslant\frac{1}{n}.$
It follows that
$\lVert u_{n}\rVert_{\mathcal{W}^{2,p}(W,\sigma^{*}g)}$ $\leqslant C_{1}\left(\lVert Lu_{n}\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}+\lVert u_{n}\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}\right)\leqslant 2C_{1}.$
Hence there exists a function $u\in\mathcal{C}^{1,\alpha}(W)$ such that $u_{n}\rightarrow u$ in the $\mathcal{C}^{1,\alpha}$-sense. Therefore,
$\lVert u\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}=1,$ $u|_{\partial W}=0\;{\rm weakly},$
$-\int_{W}4\nabla_{V}u\cdot\nabla_{V}\phi\ \text{dvol}_{\sigma^{*}g}$ $+\int_{W}4\nabla_{\sigma^{*}g}u\cdot\nabla_{\sigma^{*}g}\phi\ \text{dvol}_{\sigma^{*}g}+\int_{W}uL_{0}(\phi)\text{dvol}_{\sigma^{*}g}$
$+\int_{W}R_{g}|_{\sigma(W)}u\phi\ \text{dvol}_{\sigma^{*}g}=0,\forall\phi\in\mathcal{C}_{c}^{\infty}(W).$
Thus $Lu=0$ weakly, so by the injectivity of $L$, we get $u=0$. This contradicts $\lVert u_{n}\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}=1$.
By (11), (12),
(13) $\lVert u\rVert_{\mathcal{W}^{2,p}(W,\sigma^{*}g)}\leqslant C_{1}(1+C_{1}^{\prime})\lVert Lu\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}.$
Since $u$ is smooth, the Sobolev embedding gives
$\lVert u\rVert_{\mathcal{C}^{1,\alpha}(W)}\leqslant C_{2}\lVert u\rVert_{\mathcal{W}^{2,p}(W,\sigma^{*}g)}\leqslant C_{1}C_{2}(1+C_{1}^{\prime})\lVert F\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)},$
with $C_{2}=C_{2}(W,\sigma^{*}g,L).$ Given $\eta>0$, choose $\delta$ such that $C_{1}C_{2}(1+C_{1}^{\prime})\delta\leqslant\eta$. For an arbitrary $C>0$, we can construct $F$ in Lem. 2.2 such that $\lVert F\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}<\delta.$
We need a technical lemma to control the second derivative of the solution $u$ of (7) in the $t$ direction (see (3)).
Lemma 2.3.
For any $\eta^{\prime}>0$, we can shrink $\eta$ and $\epsilon$ in the notation of Prop. 2.1 so that the solution $u$ of (7) satisfies
(14) $\left\lvert\frac{\partial^{2}u}{\partial t^{2}}\right\rvert<\eta^{\prime}$
on $X\times\left(-\frac{\epsilon}{4},\frac{\epsilon}{4}\right).$ In particular, under the identification $W\approx W\times\{P\}$, we have this estimate on $X\times\{0\}\times\{P\}\subset M.$
Proof.
By the definitions of $\sigma$ and $g$, the induced metric $\sigma^{*}g$, and the vector field $V$ do not contain $t$ derivatives. In particular, $R_{g}$ is constant on all $t$-fibers of $M$. For fixed positive $\epsilon\ll 1$, set
$U_{1}$ $:=X\times\left(-\frac{\epsilon}{4},\frac{\epsilon}{4}\right)\subset U_{2}:=X\times\left(-\frac{\epsilon}{2},\frac{\epsilon}{2}\right)\subset W,$
$V_{1}$ $:=X\times\left(-\frac{1}{4},\frac{1}{4}\right)\subset V_{2}:=X\times\left(-\frac{1}{2},\frac{1}{2}\right).$
Applying $\frac{\partial^{2}}{\partial t^{2}}$ to both sides of (7), we have
(15) $-4\nabla_{V}\nabla_{V}\frac{\partial^{2}u}{\partial t^{2}}+4\Delta_{\sigma^{*}g}\frac{\partial^{2}u}{\partial t^{2}}+R_{g}|_{\sigma(W)}\frac{\partial^{2}u}{\partial t^{2}}=0\Rightarrow L_{\sigma^{*}g}\left(\frac{\partial^{2}u}{\partial t^{2}}\right)=0\;{\rm on}\;U_{2},$
since $F$ is independent of $t$ on $U_{2}$ by Lem. 2.2.
Fix $s>0$ such that $s-\frac{n}{2}\geqslant 2-\frac{n}{p}$, where $p$ is given in Prop. 2.1. For the metric $\sigma^{*}g$, we have the Sobolev space/$H^{s}$ basic elliptic estimate for $L=L_{\sigma^{*}g}$:
(16) $\lVert v\rVert_{H^{s}(U_{1},\sigma^{*}g)}\leqslant D_{s}\left(\lVert Lv\rVert_{H^{s-2}(U_{2},\sigma^{*}g)}+\lVert v\rVert_{\mathcal{L}^{2}(U_{2},\sigma^{*}g)}\right),\forall v\in H^{s}(U_{1},\sigma^{*}g),$
where $D_{s}=D_{s}(L,\sigma^{*}g,U_{1},U_{2})$ implicitly depends on $\epsilon.$
The problem is that $D_{s}$ blows up as $\epsilon\to 0$. As a result, we need to scale both the time direction $t$ and the space dimensions $W$ so that $U_{1},U_{2}$ become $V_{1},V_{2}$, resp., and (15) is unchanged. Therefore, we set $t^{\prime}=\epsilon^{-1}t$, $g^{\prime}=\epsilon^{-2}g.$ Then (15) becomes
(17) $-4\nabla_{V^{\prime}}\nabla_{V^{\prime}}\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}+4\Delta_{\sigma^{*}g^{\prime}}\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}+R_{g^{\prime}}|_{\sigma(W)}\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}$ $=0\Rightarrow L_{\sigma^{*}g^{\prime}}\left(\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right)=0\;{\rm on}\;V_{2},$
where $V^{\prime}=\epsilon V$ is defined by (3) for $g^{\prime}.$
Combining (16) for $\sigma^{*}g^{\prime}$ with (17), we obtain
(18) $\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{H^{s}(V_{1},\sigma^{*}g^{\prime})}$ $\leqslant D_{s}\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{\mathcal{L}^{2}(V_{2},\sigma^{*}g^{\prime})},$
where $D_{s}$ and all “$D$” constants below do not depend on $\epsilon$ or $u.$
Using the basic elliptic estimate, We have
$\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{\mathcal{L}^{2}(V_{2},\sigma^{*}g^{\prime})}$ $=\epsilon^{2-\frac{n}{2}}\left\lVert\frac{\partial^{2}u}{\partial t^{2}}\right\rVert_{\mathcal{L}^{2}(V_{2},\sigma^{*}g)}\leqslant\epsilon^{2-\frac{n}{2}}\left\lVert u\right\rVert_{H^{2}(V_{2},\sigma^{*}g)}$
(19) $\leqslant\epsilon^{2-\frac{n}{2}}D_{1}\left(\lVert L_{g}u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g)}+\left\lVert u\right\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g)}\right)$
$\leqslant\epsilon^{2-\frac{n}{2}}D_{1}\left(\lVert F\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g)}+D_{1}^{\prime}\lVert u\rVert_{\mathcal{L}^{\frac{2(n+1)}{n-1}}(W,\sigma^{*}g)}\right)$
$\leqslant D_{1}\epsilon^{2}\lVert F\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime}))}+D_{1}D_{1}^{\prime}\epsilon^{2-\frac{n}{n+1}}\lVert u\rVert_{\mathcal{L}^{\frac{2(n+1)}{n-1}}(W,\sigma^{*}g^{\prime})}.$
Since $g=h\oplus dt^{2}$, the mean curvature satisfies $h_{g}\equiv 0$. Also, since $R_{g}>0$ and $\dim M=n+1$, the relative Yamabe constant [2] $\lambda:=\lambda(M,\partial M,[g])$ for the conformal class $[g]$ on $(M,\partial M)$ is positive and satisfies
(20) $\lVert v\rVert_{\mathcal{L}^{\frac{2(n+1)}{n-1}}(M,g)}^{2}\leqslant\lambda^{-1}\left(\frac{4n}{n-1}\lVert\nabla_{g}v\rVert_{\mathcal{L}^{2}(M,g)}^{2}+\int_{M}R_{g}v^{2}{\rm dvol}_{g}\right),\forall v\in H^{1}(M,g).$
$\lambda$ is a conformal invariant, so we can replace $g$ by $g^{\prime}$ in (20).
Let $\bar{u}=u\circ\pi:M\to\mathbb{R}$ be the pullback of $u:W\to\mathbb{R}$ under the projection $\pi:M\to W$, so $\bar{u}$ is constant on the $t$-fibers. Then for $\theta^{\prime}=\epsilon^{-1}\theta,$
[…]
where $D_{2}$ satisfies $\text{dvol}_{d(\theta^{\prime})^{2}\oplus\sigma^{*}g^{\prime}}\leq D_{2}^{\frac{2(n+1)}{n-1}}\ \text{dvol}_{g^{\prime}}$. The fact that $D_{2}$ is independent of $\epsilon$ (and $u$) is the reason we scale $\theta.$
Using (20) for $g^{\prime}$, (21) becomes
${\lVert u\rVert_{\mathcal{L}^{\frac{2(n+1)}{n-1}}(W,\sigma^{*}g^{\prime})}}$
$\leqslant\epsilon^{\frac{n-1}{2(n+1)}}D_{2}\lambda^{-1/2}\left(\frac{4n}{n-1}\lVert\nabla_{g^{\prime}}\bar{u}\rVert^{2}_{\mathcal{L}^{2}(M,g^{\prime})}+\int_{M}R_{g^{\prime}}\bar{u}^{2}\text{dvol}_{g^{\prime}}\right)^{\frac{1}{2}}$
(22) $\leqslant\epsilon^{\frac{n-1}{2(n+1)}}D_{2}D_{3}\lambda^{-1/2}\left(\frac{4n}{n-1}\cdot\frac{1}{2\pi}\int_{\mathbb{S}^{1}}d\theta^{\prime}\int_{W}\lvert\nabla_{\sigma^{*}g^{\prime}}u\rvert^{2}_{\sigma^{*}g^{\prime}}\text{dvol}_{\sigma^{*}g^{\prime}}\right.$
$\qquad\left.+\frac{1}{2\pi}\int_{\mathbb{S}^{1}}d\theta^{\prime}\int_{W}R_{g^{\prime}}u^{2}\text{dvol}_{\sigma^{*}g^{\prime}}\right)^{\frac{1}{2}}$
$=D_{2}D_{3}\lambda^{-1/2}\left(\frac{4n}{n-1}\lVert\nabla_{\sigma^{*}g^{\prime}}u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime})}^{2}+\epsilon^{2}\max_{W}\lvert R_{g}\rvert\lVert u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime})}\right)^{\frac{1}{2}},$
where we have $\text{dvol}_{g^{\prime}}\leq D_{3}^{2}\text{dvol}_{d(\theta^{\prime})^{2}\oplus\sigma^{*}g^{\prime}}$. We emphasize that the introduction of $\bar{u}$ lets us use the positivity of $R_{g}$ and hence $\lambda$ on $M$ to get the estimate (2) for any function $u$ on $W$.
We also have
(23) $\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{\mathcal{C}^{0}(V_{1})}\leqslant D_{0}\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{H^{s}(V_{1},\sigma^{*}g)}=D_{0}\epsilon^{\frac{n}{2}}\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{H^{s}(V_{1},\sigma^{*}g^{\prime})}.$
The first inequality uses the Sobolev embedding theorem. For the second inequality, the zeroth order contribution to the $H^{s}$ norm contributes $\epsilon^{\frac{n}{2}}$ from the scaling of the volume form; higher order derivatives in $t$ contribute higher powers of $\epsilon,$ which can be ignored.
Applying (8), (18), (2), (2) to (23), we get
${\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{\mathcal{C}^{0}(V_{1})}\leqslant D_{0}D_{s}\epsilon^{\frac{n}{2}}\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{\mathcal{L}^{2}(V_{2},\sigma^{*}g^{\prime})}}$
$\leqslant D_{0}D_{1}D_{1}^{\prime}D_{s}\epsilon^{2-\frac{n}{n+1}+\frac{n}{2}}\lVert u\rVert_{\mathcal{L}^{\frac{2(n+1)}{n-1}}(W,\sigma^{*}g^{\prime})}+D_{0}D_{1}D_{s}\epsilon^{2+\frac{n}{2}}\lVert F\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime})}$
$\leqslant D_{0}D_{1}D_{1}^{\prime}D_{2}D_{3}D_{s}\lambda^{-1/2}\epsilon^{2-\frac{n}{n+1}+\frac{n}{2}}\left(\frac{4n}{n-1}\lVert\nabla_{\sigma^{*}g^{\prime}}u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime})}^{2}+\epsilon^{2}\max_{W}\lvert R_{g}\rvert\lVert u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime})}\right)^{\frac{1}{2}}$
$\qquad+D_{0}D_{1}D_{s}\epsilon^{2+\frac{n}{2}}\lVert F\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime})}$
$=D_{0}D_{1}D_{1}^{\prime}D_{2}D_{3}D_{s}\lambda^{-1/2}\epsilon^{2-\frac{n}{n+1}+1}\left(\frac{4n}{n-1}\lVert\nabla_{\sigma^{*}g}u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g)}^{2}+\max_{W}\lvert R_{g}\rvert\lVert u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g)}\right)^{\frac{1}{2}}$
$\qquad+\epsilon^{2}\lVert F\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g)}$
$<\bar{D}\epsilon^{2+\frac{1}{n+1}}\eta+\bar{D}^{\prime}\epsilon^{2}(C+1)\epsilon^{1/2},$
for positive constants $\bar{D},\bar{D}^{\prime}$. In the last line, $\eta$ comes from Prop. 2.1, with $C$ fixed and $\epsilon$ in this proposition to be determined; for the last term, we use $\text{supp}(F)\subset X\times[-\epsilon,\epsilon]$ in Lem. 2.2. If the estimate in this lemma holds for a fixed $C$ and $\epsilon_{0}$, then it holds for all $\epsilon<\epsilon_{0}.$ Therefore,
$\left\lVert\frac{\partial^{2}u}{\partial t^{2}}\right\rVert_{\mathcal{C}^{0}(U_{1})}=\epsilon^{-2}\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{\mathcal{C}^{0}(V_{1})}<\bar{D}\epsilon^{\frac{1}{n+1}}\eta+\bar{D}^{\prime}(C+1)\epsilon^{1/2}<\eta^{\prime},$
by fixing $\eta$ and shrinking $\epsilon$ as needed. ∎
Remark 2.1.
(i) The only places in the paper we use the existence of a PSC metric $h$ on $X\times{\mathbb{S}}^{1}$ are in (10) and in the positivity of the relative Yamabe invariant above (20). The only place we use that $X$ and hence $W$ are oriented is doing the integration by parts/Stokes’ Theorem in (2).
(ii) The solution $u$ of (7) on $W$ may not be positive. However, $u_{W}:=u+1$ is positive, provided $\eta\ll 1$. Let $u_{M}$ be the pullback of $u_{W}$ to $M$ for the projection $M\to W$, and let $u_{Y}$ be the restriction of $u_{M}$ to $Y=X\times\{0\}\times\mathbb{S}^{1}=X\times S^{1}.$ Note that $u_{W}$ satisfies
(24) $4\nabla_{V}\nabla_{V}u_{W}-4\Delta_{\imath^{*}g}u_{W}+R_{g}|_{\sigma(W)}u_{W}$ $=F+R_{g}|_{\sigma(W)}\;{\rm in}\;W,u_{W}=1\;{\rm on}\;\partial W.$
$\left\lvert\frac{\partial^{2}u_{W}}{\partial t^{2}}\right\rvert<\eta^{\prime}\;{\rm on}\;X\times(-\beta,\beta)\subset W.$
(iii) We use $e^{2\phi_{M}}$ as the conformal factor on $M$, where
(25) $e^{2\phi_{M}}:=u_{M}^{\frac{4}{n-2}},\tilde{g}:=e^{2\phi_{M}}g=u_{M}^{\frac{4}{n-2}}g\;{\rm on}\;M.$
We set $e^{2\phi_{Y}}:=u_{Y}^{\frac{4}{n-2}}$ as the conformal factors on $Y=X\times{\mathbb{S}}^{1}$. (The exponent of $u_{M}$ is not the usual one. For $X\times{\mathbb{S}}^{1}$, the exponent is standard since $n=\dim(X\times\mathbb{S}^{1})$.) Similarly, set $e^{2\phi_{W}}:=(u_{W})^{\frac{4}{n-2}}$ on $W$. As pullbacks, $\phi_{M}$ and $u_{M}$ are constant in the $\mathbb{S}^{1}$-direction:
$\frac{\partial\phi_{M}}{\partial\theta}=\frac{\partial u_{M}}{\partial\theta}=0.$
(iv) The PDE (7) is related to the conformal change
(26) $\tilde{g}=u^{\frac{4}{n-2}}g,$
for a positive, smooth function $u$ and a metric $g$ on any manifold. For $\phi$ defined by
(27) $e^{2\phi}=u^{\frac{4}{n-2}},$
it is straightforward to check by differentiating (27) that the Laplacians of these two conformal factors are related by
$u^{-\frac{n+2}{n-2}}\left(-\frac{4}{n-2}\Delta_{g}u\right)=e^{-2\phi}\left(-2\Delta_{g}\phi-(n-2)\lvert\nabla_{g}\phi\rvert^{2}\right),n\geqslant 3,$
and their gradients are related by
$e^{-2\phi}\nabla_{g}\phi=\frac{2}{n-2}u^{-\frac{n+2}{n-2}}\nabla_{g}u,\ e^{-2\phi}\lvert\nabla_{g}\phi\rvert^{2}=\left(\frac{2}{n-2}\right)^{2}u^{-\frac{2n}{n-2}}\lvert\nabla_{g}u\rvert^{2}.$
Finally, we will use the following relation between $u_{Y}$ and $\phi_{Y}$, again given by differentiating (27):
$e^{-2\phi_{Y}}\left(2(n-2)\nabla_{V}\nabla_{V}\phi_{Y}+(n-2)^{2}\nabla_{V}\phi_{Y}\nabla_{V}\phi_{Y}\right)=4u_{Y}^{-\frac{n+2}{n-2}}\nabla_{V}\nabla_{V}u_{Y}.$
(v) We have
$u_{Y}(x,P)=u_{M}(x,0,P)=u_{W}(x,0),$
$\phi_{Y}(x,P)=\phi_{M}(x,0,P)=\phi_{W}(x,0).$
As mentioned above, we need to compare the Laplacians on $(M,g)$, $(W,\sigma^{*}g)$, and $(X\times\mathbb{S}^{1},h)$.
Lemma 2.4.
In the notation of (2) and Remark 2.1(iii), we have
(28) $\Delta_{g}u_{M}(w,P)-\Delta_{\sigma^{*}g}u_{W}(w)=B_{1}(u_{W},g)(w),$
for $w\in W$ and fixed $P\in\mathbb{S}^{1}.$ Here $B_{1}(u_{W},g)=B_{1}(u_{W},g,P)$ is a globally defined first order operator on $u_{W}\in C^{\infty}(W)$; in particular, the left hand side of (28) has no derivatives in the $\mathbb{S}^{1}$-direction. Similarly, we have
(29) $\Delta_{g}u_{M}(y,0)-\Delta_{\imath^{*}g}u_{Y}(y)=\frac{\partial^{2}u_{M}}{\partial t^{2}}(y,0),$
for $y\in X\times\mathbb{S}^{1}$.
Proof.
In (28), for $\pi:M\to W$, we are computing
(30) $\Delta_{g}u_{M}-\pi^{*}(\Delta_{\sigma^{*}g}u_{W})=\Delta_{g}(u_{W}\circ\pi)-\pi^{*}(\Delta_{\sigma^{*}g}(u_{W}))=\Delta_{g}(u_{W}\circ\pi)-\pi^{*}(\Delta_{\sigma^{*}g}(u_{M}\circ\sigma))$
in a tubular neighborhood $T$ of the hypersurface $H:=W\times\{P\}\subset M$, and then restricting to $H$. By taking the level surfaces for the $g$-distance function to $H$, we can take a chart $(x^{1},\ldots,x^{n},t)$ in $M$ containing a chart $(x^{1},\ldots,x^{n-1},t)$ for $W\times\{P\}$, with $t:=x^{n+1}$ the coordinate on $[-1,1]$ and $\partial_{x^{n}}$ normal to the level surfaces. If $x^{n}(P)=0$ in these coordinates, we can write $h=h_{X,x_{n}}\oplus d(x^{n})^{2}$ and $g=h_{X,x_{n}}\oplus d(x^{n})^{2}\oplus dt^{2}$ near $W\times\{P\}$. Therefore, $\sigma^{*}g=h_{X,0}\oplus dt^{2}$ on $W\approx W\times\{P\}$, and $g^{ij}=(\sigma^{*}g)^{ij},i,j=1,\dotso,n-1$ on $W$. In summary, on $T$,
$(g_{ij})=\begin{pmatrix}g_{11}&\dotso&g_{1,n-1}&0&0\\ \vdots&\dotso&\vdots&\vdots&\vdots\\ g_{n-1,1}&\dotso&g_{n-1,n-1}&0&0\\ 0&\dotso&0&1&0\\ 0&\dotso&0&0&1\end{pmatrix},$
with $(g^{ij})$ the inverse of the whole matrix, and $(\sigma^{*}g)^{ij}$ the inverse of the minor given by deleting the $n^{\rm th}$ row and column. Then
$\Delta_{g}u_{M}=\sum_{i,j=1}^{n-1}g^{ij}\frac{\partial^{2}u_{M}}{\partial x^{i}\partial x^{j}}+\frac{\partial^{2}u_{M}}{\partial(x^{n})^{2}}+\frac{\partial^{2}u_{M}}{\partial t^{2}}-\sum_{i,j,k=1}^{n+1}g^{ij}\Gamma_{ij,g}^{k}\frac{\partial u_{M}}{\partial x^{k}}.$
The projection map $\pi:M\rightarrow W$ is $\pi(x^{1},\dotso,x^{n-1},\theta,t)=(x^{1},\dotso,x^{n-1},t)$, so $\frac{\partial u_{M}}{\partial x^{n}}=\frac{\partial(u_{W}\circ\pi)}{\partial x^{n}}=0.$ Similarly, it is easy to check that
$\frac{\partial u_{W}}{\partial x^{i}}\biggl{|}_{w}=\frac{\partial(u_{M}\circ\sigma)}{\partial x^{i}}\biggl{|}_{w}=\frac{\partial u_{M}}{\partial x^{i}}\biggl{|}_{(w,P)},i=1,\dotso,n-1,n+1.$
On $W$,
$\Delta_{\sigma^{*}g}u_{W}=\sum_{i,j=1}^{n-1}(\sigma^{*}g)^{ij}\frac{\partial^{2}u_{W}}{\partial x^{i}\partial x^{j}}+\frac{\partial^{2}u_{W}}{\partial t^{2}}-(\sigma^{*}g)^{ij}\Gamma_{ij,\sigma^{*}g}^{k}\frac{\partial u_{W}}{\partial x^{k}}.$
Since $(g^{ij})=(\sigma^{*}g)^{ij}$ on $W$, we have
$\Delta_{g}u_{M}(w,P)-\Delta_{\sigma^{*}g}u_{W}(w)$
$\qquad=\left(\sum_{i,j=1}^{n-1}g^{ij}\frac{\partial^{2}u_{M}}{\partial x^{i}\partial x^{j}}\biggl{|}_{(w,P)}+\frac{\partial^{2}u_{M}}{\partial t^{2}}\biggl{|}_{(w,P)}-\sum_{i,j,k=1}^{n+1}g^{ij}\Gamma_{ij,g}^{k}\frac{\partial u_{W}}{\partial x^{k}}\biggl{|}_{w}\right)$
$\qquad\qquad-\left(\sum_{i,j=1}^{n-1}(\sigma^{*}g)^{ij}\frac{\partial^{2}u_{M}}{\partial x^{i}\partial x^{j}}\biggl{|}_{(w,P)}+\frac{\partial^{2}u_{M}}{\partial t^{2}}\biggl{|}_{(w,P)}-\sum_{i,j,k\neq n}(\sigma^{*}g)^{ij}\Gamma_{ij,\sigma^{*}g}^{k}\frac{\partial u_{W}}{\partial x^{k}}\biggl{|}_{w}\right)$
$\qquad:=B_{1}(u_{W},g)(w).$
(29) follows from the fact that $g=\imath^{*}g\oplus dt^{2}$ with $t=x^{n+1}$, and hence all Christoffel symbols involving the $x^{n+1}$ direction vanish. ∎
Lemma 2.5.
For fixed $\eta$ in Prop. 2.1, for fixed $P\in{\mathbb{S}}^{1}$ and $u_{W}$ as in Remark 2.1(ii), there exists a constant $K_{1}=K_{1}(\eta)>0$ such that
(31) $4|B_{1}(u_{W},g,P)(w)|<K_{1},$
for all $w\in W.$ We have $K_{1}\to 0$ as $\eta\to 0.$
Proof.
$B_{1}$ is a first order operator in $u$ with smooth coefficients, and which does not contain derivatives of $u$ in the normal or ${\mathbb{S}}^{1}$ directions. Since the coefficients of $B_{1}$ are bounded on $W$ and we have $\|u\|_{\mathcal{C}^{1,\alpha}}<\eta$, the result follows. ∎
3. Proof of The Main Theorem
In this section, we prove Thm.  3.1. One direction of the ${\mathbb{S}}^{1}$-Stability Conjecture direction is trivial: if $X$ admits a PSC metric $g_{0}$, then the product metric on $X\times\mathbb{S}^{1}$ has PSC.
As a possible approach to the ${\mathbb{S}}^{1}$-stability conjecture in the other direction, if we start with $R_{h}=R_{\imath^{*}g}>0$ on $X\times{\mathbb{S}}^{1}$, then $R_{\imath^{*}g}|_{X\times\{P\}}>0$. In the notation of (2), if we somehow knew that $R_{\imath^{*}g}|_{X\times\{P\}}-R_{\tau^{*}\imath^{*}g}|_{X}<0$ by some Gauss-Codazzi manipulations, then $R_{\tau^{*}\imath^{*}g}>0$ on $X$. Unfortunately, it is very hard to estimate the second order derivatives of the metric in the Gauss-Codazzi equation. Instead, we make two modifications: (i) we introduce the extra dimension $[-1,1]$ in order to produce the elliptic operator $L^{\prime}$ on $W$ in Lem. 2.1; (ii) motivated by [9], we consider a conformal transformation of $g$ to $\tilde{g}$ as in (25). In terms of the diagram (2), the extra dimension in (i) allows us to transfer geometric information up the right hand side, i.e., from $(W,\sigma^{*}g)$ to $(M,g)$. The conformal transformation in (ii) produces $(M,\tilde{g})$ in the upper right corner, and Gauss-Codazzi moves us first to the upper left corner and then down to $(X,\tau^{*}\imath^{*}\tilde{g})$ in the lower left corner. In the end, we will prove that $\tau^{*}\imath^{*}\tilde{g}$ is a PSC metric on $X$, provided (1) holds.
Theorem 3.1.
Let $X$ be an oriented closed manifold with $\dim X=n-1\geqslant 2$. If $X\times\mathbb{S}^{1}$ admits a PSC metric $h$ such that for some $P\in{\mathbb{S}}^{1}$, we have
(32) $\angle_{h}(\mu,\partial_{\theta})<\frac{\pi}{4}$
on $X\times\{P\},$ then $X\approx X\times\{P\}$ admits a PSC metric.
Most of the proof works with a general conformal transformation $u_{W}$ of $\sigma^{*}g$ on $W$; it is only in (3) that we specify that $u_{W}$ satisfies (7).
Proof.
We collect a few facts in preparation for the proof.
  • We fix $p\in\mathbb{N},p\gg 0$, such that $\mathcal{W}^{2,p}(W,\sigma^{*}g)\hookrightarrow\mathcal{C}^{1,\alpha}(W)$ is a compact inclusion.
  • Pick $\eta\ll 1$ and $\delta=\delta(\eta)>0$ such that for the function $F=F(p,C,\delta)$ in Lem. 2.2, the solution $u_{W}$ of (24) satisfies $\frac{1}{2}<\|u_{W}\|_{\mathcal{C}^{0}}<\frac{3}{2}$; automatically, $\sup_{i,k}\|\partial_{x^{i}_{k}}u_{W}\|_{C^{0,\alpha}}\leqslant\eta$ by Prop. 2.1 and the notation above the Proposition. We further shrink $\eta$ as necessary so that $\eta^{\prime}<1/4$ in (14).
  • Fix $P\in{\mathbb{S}}^{1}.$ Then $|B_{1}(u_{W},g,P)(w)|<K_{1}=K_{1}(\eta)$ by Lem. 2.5. Recall that $K_{1}\to 0$ as $\eta\to 0$, so shrinking the choice of $\eta$ above improves this estimate.
  • Choose $C=C(P,u,\eta)\gg 1$ such that
    (33) $C>\frac{3}{2}\max_{x\in X}\left(2\left\lvert{\rm Ric}_{\imath^{*}g}(\mu,\mu)(x,P)\right\rvert+h_{\imath^{*}g}^{2}(x,P)+\lvert A_{\imath^{*}g}\rvert^{2}(x,P)\right)+K_{1}+2.$
    Here $h_{\imath^{*}g},A_{\imath^{*}g}$ are the mean curvature and second fundamental form of $i^{*}g.$
  • Using the lower bound on $\|u_{W}\|$ above and $\|u_{W}\|_{\mathcal{C}^{1,\alpha}}<\eta$ by Prop. 2.1, we can choose $\eta\ll 1$ so that for all $x\in X,$
    $K_{2}(u_{W},g)(x):=\frac{4}{n-2}\left(\frac{\lvert\nabla_{\imath^{*}g}u_{W}(x,0)\rvert^{2}_{g}}{u_{W}(x,0)}+n\frac{\nabla_{V}u_{W}(x,0)\nabla_{V}u_{W}(x,0)}{u_{W}(x,0)}\right)$
    has
    (34) $\lvert K_{2}(u_{W},g)(x)\rvert\leqslant C_{3}\eta<1,$
    where $C_{3}=C_{3}(g,P,u_{W},n)$. This estimate will be used in the last step of the proof.
  • All derivatives of $u_{M},\phi_{M},u_{Y},\phi_{Y}$ vanish in the ${\mathbb{S}}^{1}$ direction, by their definitions in Remark 2.1(ii),(iii). In particular, $\nabla_{\mu}\nabla_{\mu}\phi_{Y}=\nabla_{V}\nabla_{V}\phi_{Y}$.
We now work on the left side of (2) using the Gauss-Codazzi equation. Gauss-Codazzi and the conformal transformation law for the second fundamental form on $X\times\{P\}\subset X\times\mathbb{S}^{1}$ give
$R_{\tau^{*}\imath^{*}\tilde{g}}$ $=R_{\imath^{*}\tilde{g}}-2\text{Ric}_{\imath^{*}\tilde{g}}\left(e^{-\phi_{Y}}\mu,e^{-\phi_{Y}}\mu\right)+h_{\imath^{*}\tilde{g}}^{2}-\lvert A_{\imath^{*}\tilde{g}}\rvert^{2},$
$A_{\imath^{*}\tilde{g}}(V_{1},V_{2})$ $=e^{\phi_{Y}}A_{\imath^{*}g}(V_{1},V_{2})+e^{\phi_{Y}}\frac{\partial\phi_{Y}}{\partial\mu}g(V_{1},V_{2}),\forall V_{1},V_{2}\in T(X\times\{P\}).$
It follows that
$\lvert A_{\imath^{*}\tilde{g}}\rvert^{2}$ $=e^{-2\phi_{Y}}\left(\lvert A_{\imath^{*}g}\rvert^{2}+2nh_{\imath^{*}g}\frac{\partial\phi_{Y}}{\partial\mu}+n^{2}\left(\frac{\partial\phi_{Y}}{\partial\mu}\right)^{2}\right),$
$h_{\imath^{*}\tilde{g}}^{2}$ $=e^{-2\phi_{Y}}\left(h_{\imath^{*}g}^{2}+2n\frac{\partial\phi_{Y}}{\partial\mu}h_{\imath^{*}g}+n^{2}\left(\frac{\partial\phi_{Y}}{\partial\mu}\right)^{2}\right).$
The formulas for the conformal transformation of $\text{Ric}_{\imath^{*}\tilde{g}}$ and $R_{\imath^{*}\tilde{g}}$ on $X\times\mathbb{S}^{1}$ with conformal factor $e^{2\phi_{Y}}$ are
$\text{Ric}_{\imath^{*}\tilde{g}}\left(e^{-\phi_{Y}}\mu,e^{-\phi_{Y}}\mu\right)$ $=e^{-2\phi_{Y}}\left(\text{Ric}_{\imath^{*}g}\left(\mu,\mu\right)-(n-2)(\nabla_{\mu}\nabla_{\mu}\phi_{Y}-\nabla_{\mu}\phi_{Y}\nabla_{\mu}\phi_{Y})\right)$
(35) $\qquad-e^{-2\phi_{Y}}\left(\Delta_{\imath^{*}g}\phi_{Y}+(n-2)\lvert\nabla_{\imath^{*}g}\phi_{Y}\rvert^{2}\right)g_{nn},$
$R_{\imath^{*}\tilde{g}}$ $=e^{-2\phi_{Y}}(R_{\imath^{*}g}-2(n-1)\Delta_{i^{*}g}\phi_{Y}-(n-1)(n-2)|\nabla_{\imath^{*}g}\phi_{Y}|^{2}).$
Since $g_{nn}=1$, we have
$R_{\tau^{*}\imath^{*}\tilde{g}}$ $=e^{-2\phi_{Y}}\left(R_{\imath^{*}g}-2\text{Ric}_{\imath^{*}g}\left(\mu,\mu\right)+h_{\imath^{*}g}^{2}-\lvert A_{\imath^{*}g}\rvert^{2}\right)$
(36) $\qquad+e^{-2\phi_{Y}}\left(-2(n-1)\Delta_{\imath^{*}g}\phi_{Y}-(n-1)(n-2)\lvert\nabla_{\imath^{*}g}\phi_{Y}\rvert^{2}\right)$
$\qquad+e^{-2\phi_{Y}}\left(2(n-2)\left(\nabla_{\mu}\nabla_{\mu}\phi_{Y}-\nabla_{\mu}\phi_{Y}\nabla_{\mu}\phi_{Y}\right)+2\left(\Delta_{\imath^{*}g}\phi_{Y}+(n-2)\lvert\nabla_{\imath^{*}g}\phi_{Y}\rvert^{2}\right)\right).$
Here $R_{\tau^{*}\imath^{*}\tilde{g}}$ is evaluated at $x\in X$, and the right hand side of (3) is evaluated at $y=(x,P)\in X\times\{P\}\subset Y.$
Now we replace the second order terms in $\phi_{Y}$ in (3) with terms in $u_{Y}$. Using Remark 2.1(iv), we get
(37) $R_{\tau^{*}\imath^{*}\tilde{g}}$ $=u_{Y}^{-\frac{n+2}{n-2}}\left(-2\text{Ric}_{\imath^{*}g}\left(\mu,\mu\right)u_{Y}+h_{\imath^{*}g}^{2}u_{Y}-\lvert A_{\imath^{*}g}\rvert^{2}u_{Y}\right)$
$\qquad+u_{Y}^{-\frac{n+2}{n-2}}\left(4\nabla_{V}\nabla_{V}u_{Y}-4\Delta_{\imath^{*}g}u_{Y}+R_{\imath^{*}g}u_{Y}\right)$
$\qquad+e^{-2\phi_{Y}}\left((n-2)\lvert\nabla_{\imath^{*}g}\phi_{Y}\rvert^{2}-n(n-2)\nabla_{V}\phi_{Y}\nabla_{V}\phi_{Y}\right).$
In the next equation, we use Remark 2.1(ii), Remark 2.1(iv), Lem. 2.3, and (29). In the first equality, the left hand side is evaluated at $x\in X$, and all terms on the right hand side (including $h_{\imath^{*}g}^{2}$ and $\lvert A_{\imath^{*}g}\rvert^{2}$) are evaluated at $y=(x,P)\in X\times\{P\}$; in the next inequality, the term $\Delta_{g}u_{M}$ is evaluated at $(x,0,P)\in M$; in the last equality, $K_{2}$ is evaluated at $x.$ We obtain
$R_{\tau^{*}\imath^{*}\tilde{g}}$ $=u_{Y}^{-\frac{n+2}{n-2}}\left(-2\text{Ric}_{\imath^{*}g}\left(\mu,\mu\right)u_{Y}+h_{\imath^{*}g}^{2}u_{Y}-\lvert A_{\imath^{*}g}\rvert^{2}u_{Y}\right)$
$\qquad+u_{Y}^{-\frac{n+2}{n-2}}\left(4\nabla_{V}u_{Y}\nabla_{V}u_{Y}-4\Delta_{g}u_{M}+4\frac{\partial^{2}u_{Y}}{\partial t^{2}}+R_{\imath^{*}g}u_{Y}\right)$
$\qquad+e^{-2\phi_{Y}}\left(-(n-2)\lvert\nabla_{\imath^{*}g}\phi_{Y}\rvert^{2}-n(n-2)\nabla_{V}\phi_{Y}\nabla_{V}\phi_{Y}\right)$
(38) $\geqslant u_{Y}^{-\frac{n+2}{n-2}}\left(-2\text{Ric}_{\imath^{*}g}\left(\mu,\mu\right)u_{Y}+h_{\imath^{*}g}^{2}u_{Y}-\lvert A_{\imath^{*}g}\rvert^{2}u_{Y}\right)$
$\qquad+u_{Y}^{-\frac{n+2}{n-2}}\left(4\nabla_{V}u_{Y}\nabla_{V}u_{Y}-4\Delta_{g}u_{M}-4\eta^{\prime}+R_{\imath^{*}g}u_{Y}\right)$
$\qquad+u_{Y}^{-\frac{n+2}{n-2}}\cdot\frac{4}{n-2}\left(-\frac{\lvert\nabla_{\imath^{*}g}u_{Y}\rvert^{2}}{u_{Y}}-n\frac{\nabla_{V}u_{Y}\nabla_{V}u_{Y}}{u_{Y}}\right)$
$=u_{Y}^{-\frac{n+2}{n-2}}\left(-2\text{Ric}_{\imath^{*}g}\left(\mu,\mu\right)u_{Y}+h_{\imath^{*}g}^{2}u_{Y}-\lvert A_{\imath^{*}g}\rvert^{2}u_{Y}\right)$
$\qquad+u_{Y}^{-\frac{n+2}{n-2}}\left(4\nabla_{V}u_{Y}\nabla_{V}u_{Y}-4\Delta_{g}u_{M}+R_{\imath^{*}g}u_{Y}\right)$
$\qquad+u_{Y}^{-\frac{n+2}{n-2}}\left(-K_{2}(u_{W},g)-4\eta^{\prime}\right).$
Now we replace $\Delta_{g}$ with $\Delta_{\sigma^{*}g},$ which is a transfer of information from lower right to upper right in (2). In the first inequality in (3), we use (28); in the next equality, we use that $u_{W}(x,0)=u_{Y}(x,P)$ on $X\times\{P\}\times\{0\}$ by Remark 2.1(iv), and $R_{g}=R_{\imath^{*}g}$, since $g$ is a product metric; in the last equality, we use (7). Again, $R_{\tau^{*}\imath^{*}\tilde{g}}$ and and $K_{2}(u_{W},g)$ are evaluated at $x$; $u_{Y},h_{\imath^{*}g},$ and $|A_{\imath^{*}g}|^{2}$ are evaluated at $y=(x,P)$; $u_{W}$ and $B_{1}(u_{W},g)$ are evaluated at $w=(x,0)\in W;$ $R_{g}$ is evaluated at $(x,0,P)\in M.$ Then for $u_{W}$ solving (7), we have
$R_{\tau^{*}\imath^{*}\tilde{g}}$ $\geqslant u_{Y}^{-\frac{n+2}{n-2}}\left(-2\text{Ric}_{\imath^{*}g}\left(\mu,\mu\right)u_{Y}+h_{\imath^{*}g}^{2}u_{Y}-\lvert A_{\imath^{*}g}\rvert^{2}u_{Y}\right)$
$\qquad+u_{Y}^{-\frac{n+2}{n-2}}\left(4\nabla_{V}\nabla_{V}u_{Y}-4\Delta_{\sigma^{*}g}u_{Y}+R_{\imath^{*}g}u_{Y}\right)$
$\qquad-u_{Y}^{-\frac{n+2}{n-2}}\left(4B_{1}(u_{Y},g)\right)+u_{Y}^{-\frac{n+2}{n-2}}\left(-K_{2}(u_{W},g)-4\eta^{\prime}\right)$
(39) $=\left(u_{W}\right)^{-\frac{n+2}{n-2}}\left(-2\text{Ric}_{\imath^{*}g}\left(\mu,\mu\right)u_{W}+h_{\imath^{*}g}^{2}u_{W}-\lvert A_{\imath^{*}g}\rvert^{2}(y)u_{W}\right)$
$\qquad+\left(u_{W}\right)^{-\frac{n+2}{n-2}}\left(4\nabla_{V}\nabla_{V}u_{W}-4\Delta_{\sigma^{*}g}u_{W}+R_{g}|_{\sigma(W)}u_{W}\right)$
$\qquad-\left(u_{W}\right)^{-\frac{n+2}{n-2}}\left(4B_{1}(u_{W},g)\right)+\left(u_{W}\right)^{-\frac{n+2}{n-2}}\left(-K_{2}(u_{W},g)-4\eta^{\prime}\right)$
$=\left(u_{W}\right)^{-\frac{n+2}{n-2}}\left(-2\text{Ric}_{\imath^{*}g}\left(\mu,\mu\right)u_{W}+h_{\imath^{*}g}^{2}u_{W}-\lvert A_{\imath^{*}g}\rvert^{2}u_{W}\right)$
$\qquad+\left(u_{W}\right)^{-\frac{n+2}{n-2}}\left(F+R_{g}-4B_{1}(u_{W},g)-K_{2}(u_{W},g)-4\eta^{\prime}\right).$
By Lem. 2.2, $F=C+1$ on $X\times\{0\}$, so by (10), (31), (33), and (34),
$R_{\tau^{*}\imath^{*}\tilde{g}}$ $>u_{W}^{-\frac{n+2}{n-2}}\left(C+1-\frac{3}{2}\max_{x\in X}\left(2\left\lvert\text{Ric}_{\imath^{*}g}\left(\mu,\mu\right)(y,P)\right\rvert+h_{\imath^{*}g}^{2}(x,P)+\lvert A_{\imath^{*}g}\rvert^{2}(x,P)\right)-K_{1}-2\right)$
$>0.$
Remark 3.1.
The technical difficulty in the proof is assuring that the operator $L^{\prime}$ in Lem. 2.1 is elliptic, so that we get a smooth conformal factor $u_{W}.$ The difficult term $\nabla_{\mu}\nabla_{\mu}$ in $L^{\prime}$ appears because (i) we have to work on $M=X\times{\mathbb{S}}^{1}\times[0,1]$, as explained before Lem. 2.1, and (ii) to pass from $M$ to $X$, we have to use Gauss-Codazzi for $\tilde{g}$ twice:
$R_{\tau^{*}\imath^{*}\tilde{g}}$ $=R_{\imath^{*}\tilde{g}}-2\text{Ric}_{\imath^{*}\tilde{g}}\left(e^{-2\phi_{Y}}\mu,e^{-2\phi_{Y}}\mu\right)+h_{\imath^{*}\tilde{g}}^{2}-\lvert A_{\imath^{*}\tilde{g}}\rvert^{2}$
$=R_{\tilde{g}}-2\text{Ric}_{\tilde{g}}\left(e^{-2\phi_{M}}\nu^{\prime},e^{-2\phi_{M}}\nu^{\prime}\right)+h_{\tilde{g}}^{2}-\lvert A_{\tilde{g}}\rvert^{2}$
$\qquad-2\text{Ric}_{\imath^{*}\tilde{g}}\left(e^{-2\phi_{Y}}\mu,e^{-2\phi_{Y}}\mu\right)+h_{\imath^{*}\tilde{g}}^{2}-\lvert A_{\imath^{*}\tilde{g}}\rvert^{2}.$
The term $\nabla_{\mu}\nabla_{\mu}\phi_{Y}$ then appears in (3) from the formulas for conformal transformations from $\text{Ric}_{\imath^{*}\tilde{g}}(e^{-\phi_{Y}}\mu,e^{-\phi_{Y}}\mu)$ to $\text{Ric}_{\imath^{*}g}(\mu,\mu)$. We must use these formulas, since the theorem’s hypothesis involves the metric $h=\imath^{*}g.$
There are several known classes of counterexamples to the ${\mathbb{S}}^{1}$-stability conjecture, including odd degree hypersurfaces in $\mathbb{C}\mathbb{P}^{3}$ [8, Rmk. 1.25], and $M\ \#\ k\overline{\mathbb{C}\mathbb{P}^{2}}$ for a simply connected Kähler surface $M$ and a positive integer $k$ [6, Rmk. 5]. These are $4$-manifolds $X$ which admit no PSC metric, but such that $X\times S^{1}$ admits a PSC metric. For these and any other counterexamples, the PSC metric on $X\times S^{1}$ has large angles in the following sense:
Corollary 3.1.
If $X$ admits no PSC metric, but $X\times{\mathbb{S}}^{1}$ admits a PSC metric $h$, then for all $P\in{\mathbb{S}}^{1}$, there exists $x=x(P)$ such that
$\angle_{h}(\mu,\partial_{\theta}))_{(x,P)}\geq\frac{\pi}{4}.$
There is an analytic consequence of Thm.  3.1. Let $\lambda_{g}(X)$ be the Yamabe invariant of the closed manifold $X$ with respect to the conformal class $[g]$.
Corollary 3.2.
Let $X\times\mathbb{S}^{1}$ be an oriented closed manifold with metric $h$ satisfying (1) on some $X\times\{P\}.$ Then
$\lambda_{h}(X\times\mathbb{S}^{1})>0\Rightarrow\lambda_{\tau^{*}h}(X)>0.$
Proof.
[…]